# OnePerc: A Randomness-aware Compiler for Photonic Quantum Computing

### v1

@@ 2.2. Realistic Hardware
[…]
@@ 3. Overview
To provide a concise overview, we exemplify the compilation flow by compiling the program graph state in Fig. 4(a) onto the hardware in Fig. 4(b), which is 3 layers of that in Fig. 3(b). Indeed, while Fig. 3(b) depicts only a single RSL, the incorporation of additional layers is both allowed and necessary for larger graph states. Steps (b)$\rightarrow$(d)$\rightarrow$(c) demonstrate the online pass, while step (a)$\rightarrow$(c) illustrates the offline pass. In the online pass, fusions are conducted concurrently in a predetermined pattern (Fig. 4(d)) without individual retries of the failed ones, which eliminates the necessity for sequential operations. This results in a physical graph state with a random 2D structure, which is then reshaped to a program-agnostic 2×12$\times$ 12 × 1 lattice (Fig. 3(c)), characterized by a regular structure to enhance real-time efficiency. When the fusion success probability exceeds the percolation threshold, this reshaping process attains near-deterministic success as the RSL size increases. This eliminates the necessity for repetitive retries of the entire compilation. With this determinism, the offline pass can be employed to improve the efficiency by mapping the program graph state compactly onto the 2×12$\times$ 12 × 1 lattice (bold blue lines in Fig. 3(c)).
@@ 7.1. Experiment Setup
Fusion Success Rate #Qubits Virtual Hardware Size RSL Size
0.90 4 2x2 24x24
9 3x3 36x36
25 5x5 60x60
0.75 4 2x2 48x48
25 5x5 120x120
64 8x8 192x192
@@
Fusion Success Rate Benchmark Name OneQ #RSL OnePerc #RSL #RSL Improv. OneQ #Fusion OnePerc #Fusion #Fusion Improv.
0.90 (hyper-advanced) QAOA-4 304 84 3.62 13,990 117,664 0.12
QFT-4 3,759 174 21.59 180,634 274,155 0.66
RCA-4 3,107 237 13.11 63,814 373,646 0.17
VQE-4 56 22 2.55 1,707 33,526 0.05
QAOA-9 $>10^{6}$ 240 $>10^{3}$ $>10^{10}$ 855,354 $>10^{4}$
QFT-9 570 $>10^{3}$ 2,031,813 $>10^{3}$
RCA-9 1,017 $>10^{2}$ 3,627,950 $>10^{3}$
VQE-9 156 $>10^{3}$ 555,065 $>10^{4}$
QAOA-25 768 $>10^{3}$ 7,637,711 $>10^{3}$
QFT-25 2,418 $>10^{2}$ 24,065,102 $>10^{2}$
RCA-25 3,111 $>10^{2}$ 30,962,172 $>10^{2}$
VQE-25 705 $>10^{3}$ 7,010,656 $>10^{3}$
0.75 (practical) QAOA-4 1,708 48 35.58 119,731 169,431 0.71
QFT-4 $>10^{6}$ 210 $>10^{3}$ $>10^{10}$ 746,977 $>10^{4}$
RCA-4 $>10^{6}$ 201 $>10^{3}$ $>10^{10}$ 714,835 $>10^{4}$
VQE-4 1,017 23 44.22 25,354 96,332 0.26
QAOA-25 $>10^{6}$ 705 $>10^{3}$ $>10^{10}$ 15,781,250 >10>10> 10
QFT-25 2,271 $>10^{2}$ 50,835,771
RCA-25 3,252 $>10^{2}$ 72,795,212
VQE-25 759 $>10^{3}$ 17,292,345
QAOA-64 2,787 $>10^{2}$ 62,386,302
QFT-64 9,609 215,095,078
RCA-64 8,001 179,100,397
VQE-64 3,132 70,109,041
@@ 7.3. Scalability and Parallelism
Figure 11. Scalability and parallelism of OnePerc with 7-qubit resource states, with fusion success probability being 0.75 for (b)(c), average node size chosen as $\times 20$ 20 for (b)(c), hardware size being $\times 100$ 100 for (b) and $\times 240$ 240 for (c).

### v2

@@ 2.2. Realistic Hardware
With the advanced integrated silicon photonics, hardware components described above can operate on the scales of GHz clock rates (GHz_1, ; GHz_2, ; GHz_3, ), leading to a time scale $\sim$ 1 ns for RSG cycles. Spatial routing can be adjusted in every RSG cycle with switches, while temporal routing can be achieved by temporarily storing photonic qubits in a high-capacity quantum memory known as delay lines, which can be realized with current optical fiber technology. With a low transmission loss rate of $<5\%$ per km (fiber, ), photons can have a lifetime of around 5000 RSG cycles in the delay lines. Moreover, the size of RSL is not completely constrained by the number of RSGs, but can be extended by leveraging the tradeoff between spatial and temporal fusions (interleaving, ), with the 5000-cycle photon lifetime allowing a 5000$\times$ extension at most.
@@ 3. Overview
To provide a concise overview, we exemplify the compilation flow by compiling the program graph state in Fig. 4(a) onto the hardware in Fig. 4(b), which is 3 layers of that in Fig. 3(b). Indeed, while Fig. 3(b) depicts only a single RSL, the incorporation of additional layers is both allowed and necessary for larger graph states. Steps (b)$\rightarrow$(d)$\rightarrow$(c) demonstrate the online pass, while step (a)$\rightarrow$(c) illustrates the offline pass. In the online pass, fusions are conducted concurrently in a predetermined pattern (Fig. 4(d)) without individual retries of the failed ones, which eliminates the necessity for sequential operations. This results in a physical graph state with a random 2D structure, which is then reshaped to a program-agnostic $2\times 1$ lattice (Fig. 3(c)), characterized by a regular structure to enhance real-time efficiency. When the fusion success probability exceeds the percolation threshold, this reshaping process attains near-deterministic success as the RSL size increases. This eliminates the necessity for repetitive retries of the entire compilation. With this determinism, the offline pass can be employed to improve the efficiency by mapping the program graph state compactly onto the $2\times 1$ lattice (bold blue lines in Fig. 3(c)).
@@ 7.1. Experiment Setup
Fusion Success Rate
#Qubits
Virtual Hardware Size
RSL Size
0.90
4
2x2
24x24
9
3x3
36x36
25
5x5
60x60
0.75
4
2x2
48x48
25
5x5
120x120
64
8x8
192x192
@@
Fusion Success Rate
Benchmark Name
OneQ #RSL
OnePerc #RSL
#RSL Improv.
OneQ #Fusion
OnePerc #Fusion
#Fusion Improv.
0.90 (hyper-advanced)
QAOA-4
304
84
3.62
13,990
117,664
0.12
QFT-4
3,759
174
21.59
180,634
274,155
0.66
RCA-4
3,107
237
13.11
63,814
373,646
0.17
VQE-4
56
22
2.55
1,707
33,526
0.05
QAOA-9
$>10^{6}$
240
$>10^{3}$
$>10^{10}$
855,354
$>10^{4}$
QFT-9
570
$>10^{3}$
2,031,813
$>10^{3}$
RCA-9
1,017
$>10^{2}$
3,627,950
$>10^{3}$
VQE-9
156
$>10^{3}$
555,065
$>10^{4}$
QAOA-25
768
$>10^{3}$
7,637,711
$>10^{3}$
QFT-25
2,418
$>10^{2}$
24,065,102
$>10^{2}$
RCA-25
3,111
$>10^{2}$
30,962,172
$>10^{2}$
VQE-25
705
$>10^{3}$
7,010,656
$>10^{3}$
0.75 (practical)
QAOA-4
1,708
48
35.58
119,731
169,431
0.71
QFT-4
$>10^{6}$
210
$>10^{3}$
$>10^{10}$
746,977
$>10^{4}$
RCA-4
$>10^{6}$
201
$>10^{3}$
$>10^{10}$
714,835
$>10^{4}$
VQE-4
1,017
23
44.22
25,354
96,332
0.26
QAOA-25
$>10^{6}$
705
$>10^{3}$
$>10^{10}$
15,781,250
$>10$
QFT-25
2,271
$>10^{2}$
50,835,771
RCA-25
3,252
$>10^{2}$
72,795,212
VQE-25
759
$>10^{3}$
17,292,345
QAOA-64
2,787
$>10^{2}$
62,386,302
QFT-64
9,609
215,095,078
RCA-64
8,001
179,100,397
VQE-64
3,132
70,109,041
@@ 7.3. Scalability and Parallelism
Figure 11. Scalability and parallelism of OnePerc with 7-qubit resource states, with fusion success probability being 0.75 for (b)(c), average node size chosen as $20\times 20$ for (b)(c), hardware size being $100\times 100$ for (b) and $240\times 240$ for (c).